# ReferencesSuppReferences

### v4

@@ 3 Properties of the OLS Event Study Estimator
4 Empirical Analysis of Identification
@@
In Section 4.1 we present a simple procedure that can be used as a sensitivity analysis to determine whether relative exogeneity provides a good approximation in practice and apply it to the analysis of interest rate responses to monetary policy shocks. In Section 4.2 we discuss some identification issues in the event study regressions that involve Blue Chip forecasts about real GDP growth.
@@ 4.1 Response of Interest Rates to Monetary Policy News
We propose a simple empirical procedure to check if $\sigma_{e,t}^{2}$ is relatively large enough that $\widehat{\beta}_{\mathrm{ES}}$ has low bias. We apply our analysis to the setting of Nakamura and Steinsson (2018a) for concreteness, noting that it can be straightforwardly applied to any other event study regression. While it is possible to estimate $\mathrm{Var}\left(D_{t}\right)$ using the sample variance of $D_{t}$ in the policy sample, this could be very large even when relative exogeneity is satisfied only for a few announcements since the variance cannot be negative, as discussed above. Additionally, it is more difficult to estimate $\sigma_{\widetilde{u},t}^{2}$ since $\widetilde{u}_{t}$ is not observed and the OLS residuals may not be close in probability to the corresponding true errors $\widetilde{u}_{t}$ given that $\mathbb{E}\left(\widetilde{u}_{t}|\,D_{t}\right)\neq 0$. The key idea to our procedure is that since $\widetilde{u}_{t}$ includes macroeconomic news or factors that are present even when there is no announcement, the order of magnitude of its variance can be retrieved from the order of magnitude of the variance of $Y_{t}$ in the control sample $\mathbf{C}$. In fact, in the control sample the variance of $D_{t}$ and $\widetilde{u}_{t}$ have the same order of magnitude and given that $\left|\beta\right|<\infty$, the order of magnitude of the variance of $Y_{t}$ for $t\in\mathbf{C}$ is the same as that of $\widetilde{u}_{t}$ so long as (4.1) also holds for $t\in\mathbf{C}$. Thus, we can proxy the average variance of $\widetilde{u}_{t}$ for $t\in\mathbf{P}$ by the average variance of $Y_{t}$ for $t\in\mathbf{C}$.
@@
To analyze whether the variance of the policy shock is relatively large only for some announcement days and not for others, we compute the sample variance of the policy variable $\mathrm{\widehat{Var}}\left(D_{t}\right)$ over disjoint sub-samples. We consider the full sample from 1/1/2000 to 3/19/2014 and the sample post-2004.242424We consider the sample after 2004 because the data for 2-year forward rates, are available from 2004 onward. The sample sizes are $T=T_{P}+T_{C}=1,236$ $\left(T_{P}=106,\,T_{C}=1130\right)$ and $T=T_{P}+T_{C}=836$ $\left(T_{P}=74,\,T_{C}=762\right)$, respectively. For each sample we construct several sub-samples with different window lengths according to the rule $n_{P}=\left\lfloor T_{P}^{4/5}\right\rfloor$ and $m_{P}=\left\lfloor T_{P}/n_{P}\right\rfloor$ where $n_{P}$ is the number of observations in each window and $m_{P}$ is the number of windows in the policy sample. [^25] The available data for the control sample is up to and including 2012. Thus, in constructing the sub-samples we consider all announcement and non-announcement days until the end of 2012.262626The choice of the window length $n_{P}=\left\lfloor T_{P}^{4/5}\right\rfloor$ is optimal for nonparametric smoothing under an MSE criterion. The number of windows in the control sample is set equal to that in the policy sample, i.e., $m_{C}=m_{P}$, and we set $n_{C}=\left\lfloor T_{C}/m_{C}\right\rfloor$ so that each corresponding window in the policy and control sample brackets the same period.
@@ 4.2 Response of Blue Chip Forecasts on Output to Monetary Policy News
$BCrev_{t}$ $=c+\beta D_{t}+\widetilde{u}_{t},$ (4.2)
@@ 5 Conclusions
We establish nonparametric conditions for identification of casual effects in high-frequency event studies. We show that identification can be achieved via a separability condition on the policy shock from the other variables present in the window, and relative exogeneity which refers to the variance of the policy shock being an order of magnitude larger than that of the other variables. Under these conditions we establish the causal meaning of the event study estimand, the consistency and asymptotic distribution of the event study estimator and its robustness to nonlinearities. We propose a simple procedure that can be used to assess relative exogeneity as an approximation and apply it to Nakamura and Steinsson’s (2018a) analysis on the real effects of monetary policy.

### v5

@@ 3 Properties of the OLS Event Study Estimator
as $T_{P}\rightarrow\infty$, where $\Delta(T_{P})\equiv\int_{0}^{1}\mathbb{E}[Y_{\lfloor T_{P}s\rfloor}^{*}(\underline{d})(D_{\lfloor T_{P}s\rfloor}^{*}-\mathbb{E}(D_{\lfloor T_{P}s\rfloor}^{*}))]\mathrm{d}s\rightarrow\Delta$ and
$\mathrm{aBias}\left(\widehat{\beta}_{\mathrm{ES}}\right)$ $=c\cdot\lim_{T_{P}\rightarrow\infty}\int_{0}^{1}\mathbb{E}\left[\varphi_{Y,u}\left(Z_{\lfloor T_{P}s\rfloor},\,u_{\lfloor T_{P}s\rfloor},\,\left\lfloor T_{P}s\right\rfloor\right)\right.$
×$\quad\times\left.\left(g_{D,u}\left(Z_{\lfloor T_{P}s\rfloor},\,u_{\lfloor T_{P}s\rfloor},\,\left\lfloor T_{P}s\right\rfloor\right)-\mathbb{E}\left(g_{D,u}\left(Z_{\lfloor T_{P}s\rfloor},\,u_{\lfloor T_{P}s\rfloor},\,\left\lfloor T_{P}s\right\rfloor\right)\right)\right)\right]\mathrm{d}s.$
Clearly, if relative exogeneity is “strong enough” in the sense that $c=0$, $\mathrm{aBias}(\widehat{\beta}_{\mathrm{ES}})=0$ and $\widehat{\beta}_{\mathrm{ES}}$ is asymptotically unbiased. In addition, $\widehat{\beta}_{\mathrm{ES}}$ is super-consistent, with a faster than standard $\sqrt{T_{P}}$ rate of convergence since $\sigma_{D}\rightarrow\infty$ under relative exogeneity. Under the conditions of the corollary, it is also possible to bias-correct $\widehat{\beta}_{\mathrm{ES}}$ in case $c\neq 0$ under the “sharp null” of zero MCEs at all time periods, providing a further inference refinement. Specifically, if $Y_{t}(d)$ does not depend upon $d$, Assumption 1 implies $Y_{t}=\varphi_{Y,u}(Z_{t},\,u_{t},\,t)$ so that
$\mathbb{E}\left[Y_{t}\left(D_{t}-\mathbb{E}(D_{t})\right)\right]=\mathbb{E}\left[\varphi_{Y,u}(Z_{t},\,u_{t},\,t)\left(g_{D,u}\left(Z_{t},\,u_{t},\,t\right)-\mathbb{E}\left(g_{D,u}(Z_{t},\,u_{t},\,t)\right)\right)\right],$
by Assumptions 2-3. Therefore, $\mathrm{aBias}(\widehat{\beta}_{\mathrm{ES}})$ can be consistently estimated by
$\sqrt{\frac{T_{P}}{T_{P}^{-1}\sum_{t=1}^{T_{P}}\left(D_{t}-\overline{D}\right)^{2}}}T_{P}^{-1}\sum_{t=1}^{T_{P}}Y_{t}\left(D_{t}-\overline{D}\right)$
under Assumption 8. In the following section, we discuss bias-aware inference that does not rely upon the imposition of the “sharp null”.
4 Bias Bounds, Worst-Case Coverage and Bias-Aware Inference
The bias in the high-frequency event study estimator vanishes in the limit under relative exogeneity. We consider methods to quantify how this bias can affect the worst-case properties of the estimator using a bound on the bias, and we propose a bias-aware critical value that accounts for the potential asymptotic bias of the estimator in addition to its variance.
4.1 Worst-Case Bias
We begin with deriving the worst-case (unscaled) bias $\Delta$ from Theorem 3. The following proposition produces a useful expression for this.
Proposition 1.
Under Assumptions 1-3 and 5-8,
[…]
where $\overline{\varphi}_{Y,u}\left(s\right)=\varphi_{Y,u}(Z_{\left\lfloor T_{P}s\right\rfloor},\,u_{\left\lfloor T_{P}s\right\rfloor},\,\left\lfloor T_{P}s\right\rfloor)$, $\overline{g}_{D,u}\left(s\right)=g_{D,u}(Z_{\left\lfloor T_{P}s\right\rfloor},\,u_{\left\lfloor T_{P}s\right\rfloor},\,\left\lfloor T_{P}s\right\rfloor)$ and $\rho_{Zu,T_{P}}\left(s\right)=\mathrm{Corr}(\varphi_{Y,u}(Z_{\left\lfloor T_{P}s\right\rfloor},\,u_{\left\lfloor T_{P}s\right\rfloor},\,\left\lfloor T_{P}s\right\rfloor),\,g_{D,u}(Z_{\left\lfloor T_{P}s\right\rfloor},\,u_{\left\lfloor T_{P}s\right\rfloor},\,\left\lfloor T_{P}s\right\rfloor))$ is the correlation coefficient between the two terms in the parentheses.
[…]
$|\Delta|\leq\frac{\rho_{\mathrm{max}}}{\sigma_{D}^{2}}\lim_{T_{P}\rightarrow\infty}\int_{0}^{1}\sqrt{\mathrm{Var}\left(\overline{\varphi}_{Y,u}\left(s\right)\right)}\sqrt{\mathrm{Var}\left(\overline{g}_{D,u}\left(s\right)\right)}\mathrm{d}s.$
With this result in hand, we can immediately derive the worst-case scaled bias under the same strengthening of relative exogeneity as that in Corollary 1: under Assumptions 1-8 and $\sqrt{T_{p}}/\sigma_{D}\rightarrow c\in[0,\infty)$ as $T_{P}\rightarrow\infty$,
$\sup_{|\rho_{Zu,T_{P}}\left(s\right)|\leq\rho_{\mathrm{max}}\,\forall s\in[0,1]}$ $\sqrt{T_{P}\sigma_{D}^{2}}\left|\mathbb{E}\left(\widehat{\beta}_{\mathrm{ES}}\right)-\left(\beta_{\mathrm{ES}}-\Delta(T_{P})\right)\right|$ (4.2)
$\rightarrow c\rho_{\mathrm{max}}\lim_{T_{P}\rightarrow\infty}\int_{0}^{1}\sqrt{\mathrm{Var}\left(\overline{\varphi}_{Y,u}\left(s\right)\right)}\sqrt{\mathrm{Var}\left(\overline{g}_{D,u}\left(s\right)\right)}\mathrm{d}s.$
The worst-case bias depends on $c\rho_{\mathrm{max}}$ and on the variances of components that are functions of the shock $u_{t}$ and omitted variables $Z_{t}$.
4.2 Worst-Case Asymptotic Coverage
We use the worst-case scaled bias to study the worst-case asymptotic coverage of the standard confidence intervals:
$\mathrm{CI}\left(\widehat{\beta}_{\mathrm{ES}}\right)$ $=\left[\widehat{\beta}_{\mathrm{ES}}\pm z_{1-a/2}\sqrt{J/(T_{P}\sigma_{D}^{2})}\right],$
where $z_{1-a/2}$ is the $21-a/$ 2 quantile of the standard normal distribution. The asymptotic variance $J$ can be estimated consistently as suggested in Section 3, without affecting the results.
Corollary 2.
Let Assumptions 1-6 and 8-9 hold, $\sqrt{T_{P}}/\sigma_{D}\rightarrow c\in\left[0,\,\infty\right)$ and $\mathcal{Z}\sim\mathscr{N}(0,1)$. Then:
$\inf_{|\rho_{Zu,T_{P}}\left(s\right)|\leq\rho_{\mathrm{max}}\,\forall s\in[0,1]}$ $\mathbb{P}\left(\left(\beta_{\mathrm{ES}}-\Delta(T_{P})\right)\in\mathrm{CI}\left(\widehat{\beta}_{\mathrm{ES}}\right)\right)$
$\rightarrow\mathbb{P}\left(\left|\mathcal{Z}+\frac{c\rho_{\mathrm{max}}}{\sqrt{J}}\lim_{T_{P}\rightarrow\infty}\int_{0}^{1}\sqrt{\mathrm{Var}\left(\overline{\varphi}_{Y,u}\left(s\right)\right)}\sqrt{\mathrm{Var}\left(\overline{g}_{D,u}\left(s\right)\right)}\mathrm{d}s\right|\leq z_{1-a/2}\right).$
We can study how the worst-case asymptotic coverage varies as $c\rho_{\mathrm{max}}$ and the ratio involving $\mathrm{Var}(\overline{\varphi}_{Y,u}\left(s\right))$, $\mathrm{Var}(\overline{g}_{D,u}\left(s\right))$ and $\sqrt{J}$ change. The asymptotic variance $J$ can be estimated consistently. To estimate $\mathrm{Var}(\overline{\varphi}_{Y,u}\left(s\right))$ and $\mathrm{Var}(\overline{g}_{D,u}\left(s\right))$ we assume stationarity, under which these expressions reduce to $\mathrm{Var}(\overline{\varphi}_{Y,u})$ and $\mathrm{Var}(\overline{g}_{D,u})$, respectively. Since $\overline{\varphi}_{Y,u}$ and $\overline{g}_{D,u}$ are functions of the shock $u_{t}$ and of omitted variables $Z_{t}$—both unobserved factors that are also present in non-announcement windows—we propose estimating $\mathrm{Var}(\overline{\varphi}_{Y,u})$ and $\mathrm{Var}(\overline{g}_{D,u})$ using the sample variance of the outcome and policy variables in the control sample. Specifically, we use $\widehat{\sigma}_{Y,C}^{2}$ for the outcome variable and $\widehat{\sigma}_{D,C}^{2}$ for the policy variable, where the control sample includes time windows on days without announcements. Under stationarity, it is reasonable to assume that the variance of $Y_{t}$ on control days provides an upper bound for $\mathrm{Var}(\overline{\varphi}_{Y,u})$, and similarly, the variance of $D_{t}$ on control days provides an upper bound for $\mathrm{Var}(\overline{g}_{D,u})$. See Section 5 for further details and for an application of Corollary 2 to the setting in Nakamura and Steinsson (2018a).
4.3 Bias-Aware Inference
An alternative way to deal with a vanishing bias is to adjust the critical value upward to compensate for this bias, as suggested by armstrong/kolesar:2021. Using the bound from Proposition 1, we define the bias-aware confidence interval as
$\mathrm{CI}_{\mathrm{BA}}\left(\widehat{\beta}_{\mathrm{ES}},\,c\rho_{\mathrm{max}}\right)$ $=\left[\widehat{\beta}_{\mathrm{ES}}\pm\mathrm{cv}_{1-a/2}\left(c\rho_{\mathrm{max}}\frac{\lim_{T_{P}\rightarrow\infty}\int_{0}^{1}\sqrt{\mathrm{Var}\left(\overline{\varphi}_{Y,u}\left(s\right)\right)}\sqrt{\mathrm{Var}\left(\overline{g}_{D,u}\left(s\right)\right)}\mathrm{d}s}{\sqrt{J}}\right)\sqrt{J/(T\sigma_{D}^{2})}\right],$
where $\mathrm{cv}_{1-a/2}\left(\mathcal{B}\right)$ is the bias-aware critical value defined as the number such that $\mathbb{P}(|\mathcal{Z}+\mathcal{B}|\leq\mathrm{cv}_{1-a/2}\left(\mathcal{B}\right))=1-a$. By construction the bias-aware confidence interval has correct asymptotic coverage probability but it can be conservative.
Corollary 3.
Let Assumptions 1-6 and 8-9 hold and $\sqrt{T_{P}}/\sigma_{D}\rightarrow c\in\left[0,\,\infty\right)$. We have:
$\underset{T_{P}\rightarrow\infty}{\lim}\inf_{|\rho_{Zu,T_{P}}\left(s\right)|\leq\rho_{\mathrm{max}}\,\forall s\in[0,1]}\mathbb{P}\left(\left(\beta_{\mathrm{ES}}-\Delta(T_{P})\right)\in\mathrm{CI}_{\mathrm{BA}}\left(\widehat{\beta}_{\mathrm{ES}},\,c\rho_{\mathrm{max}}\right)\right)$ $=1-a.$
5 Empirical Analysis of Identification and Inference
@@
In Section 5.1 we present a simple procedure that can be used as a sensitivity analysis to determine whether relative exogeneity provides a good approximation in practice and apply it to the analysis of interest rate responses to monetary policy shocks. In Section 5.2 we examine the worst-case asymptotic coverage of the standard confidence interval and compare the length of the bias-aware confidence interval to that of the conventional one. In Section 5.3 we discuss some identification issues in the event study regressions that involve Blue Chip forecasts about real GDP growth.
@@ 5.1 Response of Interest Rates to Monetary Policy News
We propose a simple empirical procedure to check if $\sigma_{e,t}^{2}$ is relatively large enough that $\widehat{\beta}_{\mathrm{ES}}$ has low bias. We apply our analysis to the setting of Nakamura and Steinsson (2018a) for concreteness, noting that it can be straightforwardly applied to any other event study regression. [^19] It can also be applied in settings where the researcher believes there is a specific nonlinear relationship between the economic variables. In such cases, one would need to consider the appropriate nonlinear regression model in place of (5.1). While it is possible to estimate $\mathrm{Var}\left(D_{t}\right)$ using the sample variance of $D_{t}$ in the policy sample, this could be very large even when relative exogeneity is satisfied only for a few announcements since the variance cannot be negative, as discussed above. Additionally, it is more difficult to estimate $\sigma_{\widetilde{u},t}^{2}$ since $\widetilde{u}_{t}$ is not observed and the OLS residuals may not be close in probability to the corresponding true errors $\widetilde{u}_{t}$ given that $\mathbb{E}\left(\widetilde{u}_{t}|\,D_{t}\right)\neq 0$. The key idea to our procedure is that since $\widetilde{u}_{t}$ includes macroeconomic news or factors that are present even when there is no announcement, the order of magnitude of its variance can be retrieved from the order of magnitude of the variance of $Y_{t}$ in the control sample $\mathbf{C}$. In fact, in the control sample the variance of $D_{t}$ and $\widetilde{u}_{t}$ have the same order of magnitude and given that $\left|\beta\right|<\infty$, the order of magnitude of the variance of $Y_{t}$ for $t\in\mathbf{C}$ is the same as that of $\widetilde{u}_{t}$ so long as (5.1) also holds for $t\in\mathbf{C}$. Thus, we can proxy the average variance of $\widetilde{u}_{t}$ for $t\in\mathbf{P}$ by the average variance of $Y_{t}$ for $t\in\mathbf{C}$.
@@
To analyze whether the variance of the policy shock is relatively large only for some announcement days and not for others, we compute the sample variance of the policy variable $\mathrm{\widehat{Var}}\left(D_{t}\right)$ over disjoint sub-samples. We consider the full sample from 1/1/2000 to 3/19/2014 and the sample post-2004.252525We consider the sample after 2004 because the data for 2-year forward rates, are available from 2004 onward. The sample sizes are $T=T_{P}+T_{C}=1,236$ $\left(T_{P}=106,\,T_{C}=1130\right)$ and $T=T_{P}+T_{C}=836$ $\left(T_{P}=74,\,T_{C}=762\right)$, respectively. For each sample we construct several sub-samples with different window lengths according to the rule $n_{P}=\left\lfloor T_{P}^{4/5}\right\rfloor$ and $m_{P}=\left\lfloor T_{P}/n_{P}\right\rfloor$ where $n_{P}$ is the number of observations in each window and $m_{P}$ is the number of windows in the policy sample. [^26] The available data for the control sample is up to and including 2012. Thus, in constructing the sub-samples we consider all announcement and non-announcement days until the end of 2012.272727The choice of the window length $n_{P}=\left\lfloor T_{P}^{4/5}\right\rfloor$ is optimal for nonparametric smoothing under an MSE criterion. The number of windows in the control sample is set equal to that in the policy sample, i.e., $m_{C}=m_{P}$, and we set $n_{C}=\left\lfloor T_{C}/m_{C}\right\rfloor$ so that each corresponding window in the policy and control sample brackets the same period.
@@
5.2 Worst-Case Asymptotic Coverage and Bias-Aware Confidence Intervals
We consider the high-frequency event study regression of the 2-Year real forward rates on 30-minute monetary policy surprises. We first use the bound on the bias from Proposition 1 to assess the worst-case asymptotic coverage of standard confidence intervals for the causal effect of policy surprises on real forward rates.
Table 4: Worst-case asymptotic coverage
$a=0.90$ $\sqrt{\mathrm{Var}\left(\overline{\varphi}_{Y,u}\right)\mathrm{Var}\left(\overline{g}_{D,u}\right)/J}$
$M$ 0 0.00001 0.0001 0.001 0.01 0.1 1
0.0001 0.900 0.900 0.900 0.900 0.900 0.900 0.900
0.1 0.900 0.900 0.900 0.900 0.900 0.900 0.898
0.5 0.900 0.900 0.900 0.900 0.900 0.900 0.858
1 0.900 0.900 0.900 0.900 0.900 0.898 0.737
10 0.900 0.900 0.900 0.900 0.898 0.737 0.000
100 0.900 0.900 0.900 0.898 0.737 0.000 0.000
300 0.900 0.900 0.900 0.885 0.088 0.000 0.000
1000 0.900 0.900 0.898 0.737 0.000 0.000 0.000
Worst-case asymptotic coverage probability of the 90% confidence interval. The outcome variable is the 2-Year U.S. Treasury instantaneous real forward rate and the policy variable is the 30-minute policy news surprise.
Assuming stationarity, Table 4 reports the worst-case asymptotic coverage probability of the confidence interval $\mathrm{CI}(\widehat{\beta}_{\mathrm{ES}})$ across values of $M=c\rho_{\mathrm{max}}$ and of the variance ratio $\sqrt{\mathrm{Var}\left(\overline{\varphi}_{Y,u}\right)}\times$ $\sqrt{\mathrm{Var}\left(\overline{g}_{D,u}\right)/J}$. As $M$ or the variance ratio increases, the coverage probability deteriorates. For the regression under consideration with $\rho_{\mathrm{max}}=1$, we have $M\approx\sqrt{T_{P}}/\widehat{\sigma}_{D,P}=303$ and variance ratio $\widehat{\sigma}_{Y,C}\widehat{\sigma}_{D,C}/\sqrt{\widehat{J}}=0.00013,$ which corresponds to a coverage probability of 0.8998. This suggests that the asymptotic bias in this high-frequency regression is small enough that it does not significantly distort the coverage probability of the conventional confidence interval, even under a worst-case scenario. It would seem to take an unrealistically large bias to reduce the worst-case coverage probability significantly below the nominal level in this application.
Next, we analyze the relative length of the bias-aware confidence interval compared to the conventional confidence interval in the high-frequency event study regression of 2-Year real forward rates on 30-minute policy surprises. Table 5 shows that the bias-aware and conventional confidence intervals often have the same length. The bias-aware interval is wider only for a few combinations of $M$ and the variance ratio. It would take an unrealistically large asymptotic bias for the bias-aware confidence interval to be significantly wider than the conventional one. For the values $M=\sqrt{T_{P}}/\widehat{\sigma}_{D,P}=303$ and variance ratio $\widehat{\sigma}_{Y,C}\widehat{\sigma}_{D,C}/\sqrt{\widehat{J}}=0.00013,$ which are relevant in this application, the two confidence intervals have the same length.
Table 5: Relative length of $\mathrm{CI}_{\mathrm{BA}}$ versus $\mathrm{CI}$
$\sqrt{\mathrm{Var}\left(\overline{\varphi}_{Y,u}\right)\mathrm{Var}\left(\overline{g}_{D,u}\right)/J}$
$M$ 0 0.00001 0.0001 0.001 0.01 0.1 1
0.0001 1.000 1.000 1.000 1.000 1.000 1.000 1.000
0.1 1.000 1.000 1.000 1.000 1.000 1.000 1.005
0.5 1.000 1.000 1.000 1.000 1.000 1.001 1.118
1 1.000 1.000 1.000 1.000 1.000 1.005 1.389
10 1.000 1.000 1.000 1.000 1.005 1.389 6.869
100 1.000 1.000 1.000 1.005 1.389 6.869 61.581
300 1.000 1.000 1.000 1.044 2.603 19.027 183.161
1000 1.000 1.000 1.005 1.389 6.869 61.581 608.693
Relative length of bias-aware confidence interval versus standard confidence interval. Significance level is $a=0.10$. The outcome variable is the 2-Year U.S. Treasury instantaneous real forward rate and the policy variable is the 30-minute policy news surprise.
@@ 5.3 Response of Blue Chip Forecasts on Output to Monetary Policy News
$BCrev_{t}$ $=\beta_{0}+\beta D_{t}+\widetilde{u}_{t},$ (5.2)
@@ 6 Conclusions
We establish nonparametric conditions for identification of casual effects in high-frequency event studies. We show that identification can be achieved via a separability condition on the policy shock from the other variables present in the window, and relative exogeneity which refers to the variance of the policy shock being an order of magnitude larger than that of the other variables. Under these conditions we establish the causal meaning of the event study estimand, the super-consistency and asymptotic distribution of the event study estimator and its robustness to nonlinearities. We provide bounds on the bias and use them to study the worst-case coverage properties of standard confidence intervals and to construct bias-aware inference procedures. We propose a simple procedure that can be used to assess relative exogeneity as an approximation and apply it to Nakamura and Steinsson’s (2018a) analysis on the real effects of monetary policy.
Supplemental Appendix